# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ Appendix G Auxiliary lemmas
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.1) and (A.2). Since $\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\leq\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mu_{j}$, (G.8) and (G.9) follow from Lemma B.7 in Kock and Preinerstorfer (2025) applied to $X_{1,j}$ with $\varepsilon=\varepsilon_{n}$$a=c^{-1}\varepsilon_{n}$ and recalling that $\sigma_{m,j}\leq\sigma_{m}$ (noting that our Assumption 2.1 implies the assumptions there) such that
$\leq$ $2\sigma_{m}(\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1-\frac{1}{m}}+\sigma_{m}\mathinner{\Bigl(1+\mathinner{\Bigl[\frac{\varepsilon_{n}+c^{-1}\varepsilon_{n}}{1-\varepsilon_{n}-c^{-1}\varepsilon_{n}}\Bigr]}^{\frac{1}{m}}\Bigr)}(\varepsilon_{n}+c^{-1}\varepsilon_{n})^{1-\frac{1}{m}}$
$\leq$ $2\sigma_{m}\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{1}{m}}\varepsilon_{n}^{1-\frac{1}{m}}+\sigma_{m}\mathinner{\biggl(1+\mathinner{\Bigl[\frac{c+1}{c-1}\Bigr]}^{\frac{1}{m}}\biggr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{1}{m}}\varepsilon_{n}^{1-\frac{1}{m}}$
$=$ $\sigma_{m}\mathinner{\biggl[2\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{1}{m}}+\mathinner{\biggl(1+\mathinner{\Bigl[\frac{c+1}{c-1}\Bigr]}^{\frac{1}{m}}\biggr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{1}{m}}\biggr]}\varepsilon_{n}^{1-\frac{1}{m}},$
where the second inequality used that $(0,1)\ni x\mapsto x/(1-x)$ is strictly increasing such that $\varepsilon_{n}+{c}^{-1}\varepsilon_{n}\leq 0.5+c^{-1}0.5$ implies
$\frac{\varepsilon_{n}+{c}^{-1}\varepsilon_{c}}{1-\varepsilon_{n}-{c}^{-1}\varepsilon_{n}}\leq\frac{0.5+c^{-1}0.5}{1-0.5-c^{-1}0.5}=\frac{c+1}{c-1}.$
The following lemma, which is Lemma E.1 in Kock and Preinerstorfer (2025), is an analogue to Lemma G.2, replacing $\varepsilon_{n}$ in (5) with $\varepsilon_{n}^{\prime}$ in (9).
@@
Fix $c\in(1,\infty)$$n\in\mathbb{N}$, and $\delta\in(0,1)$. Let $Z_{1},\ldots,Z_{n}$ be i.i.d. real random variables and suppose that $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$ satisfy that
@@
If $\varepsilon_{n}^{\prime}\in(0,1/2)$ where
$\varepsilon_{n}^{\prime}=c\overline{\eta}_{n}+c\sqrt{\frac{\log(6/\delta)}{2n}},$
each of (G.10)–(G.13) below holds with probability at least $1-\frac{\delta}{6}$:
$\tilde{Z}_{\lceil\varepsilon^{\prime}_{n}n\rceil}^{*}$ $\geq$ $Q_{\varepsilon^{\prime}_{n}-c^{-1}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.10)
$\tilde{Z}_{\lceil(1-\varepsilon^{\prime}_{n})n\rceil}^{*}$ $\geq$ $Q_{1-\varepsilon^{\prime}_{n}-c^{-1}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.11)
$\tilde{Z}_{\lfloor\varepsilon^{\prime}_{n}n\rfloor+1}^{*}$ $\leq$ $Q_{\varepsilon^{\prime}_{n}+c^{-1}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.12)
$\tilde{Z}_{\lfloor(1-\varepsilon^{\prime}_{n})n\rfloor+1}^{*}$ $\leq$ $Q_{1-\varepsilon^{\prime}_{n}+c^{-1}\varepsilon^{\prime}_{n}}(Z_{1}).$ (G.13)

### v4

@@ Appendix G Auxiliary lemmas
The proof is almost identical to the one of Lemma G.2, but we now apply Lemma B.3 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{d^{2}n}$ and $\epsilon=\varepsilon_{n}^{\prime}$. ∎
The following lemma, which follows from Lemma B.5 in Kock and Preinerstorfer (2025), is an analogue to Lemma G.3, replacing $\varepsilon_{n}$ in (4) with $\varepsilon_{n}^{\prime}$ in (7).
@@
Let $Z_{1},\ldots,Z_{n}$ be i.i.d. real random variables and suppose that the random variables $\tilde{Z}_{1},\ldots,\tilde{Z}_{n}$ satisfy
@@
Suppose that $\varepsilon_{n}^{\prime}$ is as defined in (7) and satisfies (9). Then, for $c_{1,n}^{\prime}$ as in (D.1) and $c_{2,n}^{\prime}$ as in (D.2), we have
$0<\varepsilon_{n}^{\prime}\min(c_{1,n}^{\prime},c_{2,n}^{\prime})\leq\varepsilon_{n}^{\prime}(c_{1,n}^{\prime}+c_{2,n}^{\prime})<1,$
and each of (G.11)–(G.14) below holds with probability at least $1-\frac{1}{d^{2}n}$:
$\tilde{Z}_{\lceil\varepsilon^{\prime}_{n}n\rceil}^{*}$ $\geq$ $Q_{c_{1,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.11)
$\tilde{Z}_{\lceil(1-\varepsilon^{\prime}_{n})n\rceil}^{*}$ $\geq$ $Q_{1-c_{2,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.12)
$\tilde{Z}_{\lfloor\varepsilon^{\prime}_{n}n\rfloor+1}^{*}$ $\leq$ $Q_{c_{2,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1});$ (G.13)
$\tilde{Z}_{\lfloor(1-\varepsilon^{\prime}_{n})n\rfloor+1}^{*}$ $\leq$ $Q_{1-c_{1,n}^{\prime}\varepsilon^{\prime}_{n}}(Z_{1}).$ (G.14)
Proof.
The first statement in the present lemma follows from (9) along with (G.6) of Lemma G.6. For the remaining statements, apply Lemma B.5 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{d^{2}n}$ (which is strictly less than one since we assume throughout that $d\geq 2$ and $n\geq 4$, cf. the sentence before Theorem 2.1), $\epsilon=\varepsilon_{n}^{\prime}$, and $\eta$ there being $\overline{\eta}_{n}$. To this end, note that
  1. 1.
    $c_{1}$, and $c_{2}$ there equal $c^{\prime}_{1,n}$ and $c^{\prime}_{2,n}$, respectively, for $\delta=\frac{6}{d^{2}n}$,
  2. 2.
    that $\varepsilon^{\prime}_{n}\geq\lambda^{\prime}_{1}\overline{\eta}_{n}$ by definition and $\varepsilon^{\prime}_{n}c^{\prime}_{2,n}<1$ (which has already been verified),
  3. 3.
    $\varepsilon^{\prime}_{n}<1/2$ follows from (9).